Johnson Industries finances its projects with 40 percent debt, 10 percent preferred stock, and 50 percent common stock.

·

The company can issue bonds at a yield to maturity of 7.1 percent.

·

The cost of preferred stock is 9 percent.

·

The company's common stock currently sells for $35 a share.

·

The company's dividend has just paid $2.00 a share (D0 = $2.00), and is expected to grow at a constant rate of 5 percent per year.

·

Assume that the flotation cost on debt and preferred stock is zero, and no new stock will be issued.

·

The company's tax rate is 30 percent


What is the company's weighted average cost of capital (WACC)? Express your answer in percentage (without the % sign) and round it to two decimal places.

.
Answer #1

The WACC of this company is :

Debt = 40%

Preference stock = 10%

Equity = 50%

Yield to maturity = 7.1%

Cost of prefernce stock = 9%

Cost of common stock =

Re = D1/ PO + G

=$2 *1.05/ $35 + 0.05

=11%

So, the WACC is :

weight of debt * after tax cost of debt + weight of prefernce stock*cost of preference stock + weight of equity *cost of equity

= 0.4* 0.071* (1- 0.3)  + 0.1*0.09 + 0.5*0.11

=0.0199+ 0.009 + 0.055

= 8.39 ( rounded off to two decimal places)